Ben is 3 years older than his sisyer and a third of the age of his mother.His father is 4 years older than his mother and four times the age of his sister. How old is Ben?
Answer by jorel1380(3719)

Let Ben's age be b. Then his sister is b-3, and his mother is 3b. This makes his father 3b+4. So:
3b+4=4(b-3)
3b+4=4b-12
b=16
Ben is 16. ☺☺☺☺